# Uncertainty-Aware PCA for Arbitrarily Distributed Data Modeled by Gaussian Mixture Models

### v1

@@ 17 Compute eigendecomposition of $\Sigma_{\mathrm{UA}}$
Dataset
Size
Dim.
Labels
GMM
Components
KL(PCA-KDE,
wGMM-UAPCA)
KL(PCA-KDE,
UAPCA)
SW2(PCA-KDE,
wGMM-UAPCA)
SW2(PCA-KDE,
UAPCA)
bank 2058 63 2
16, 8
4.0804e-04 8.1839e-04 1.2428e-02 1.9343e-02
cifar10 3249 1024 10
1, 14, 1, 1, 1
1, 1, 13, 12 14
3.6982e-03 3.1093e-03 9.6107e-03 1.0517e-02
cnae9 1079 856 9
2, 1, 1, 1, 2
2, 1, 1, 1
4.3181e-04 4.4680e-04 8.9226e-03 8.8375e-03
coil20 1439 400 20
3, 3, 3, 3, 3
3, 3, 2, 3, 2
3, 2, 3, 3, 2
1, 2, 3, 3, 3
3.9615e-03 2.6747e-03 8.6938e-03 1.2066e-02
epileptic 5749 178 5
2, 3, 2, 1, 1
3.7940e-05 6.0040e-05 5.2632e-03 6.1744e-03
fashion_mnist 2999 784 10
1, 2, 12, 10, 11
2, 1, 2, 10, 2
8.4062e-03 9.9694e-03 8.8474e-03 1.2469e-02
fmd 996 1536 10
4, 4, 4, 4, 4
4, 5, 4, 3, 5
1.5755e-03 9.4136e-04 1.4576e-02 1.2633e-02
har 734 561 6
5, 5, 4, 4, 4, 5
3.0825e-03 2.2314e-03 8.5085e-03 1.1451e-02
hatespeech 3220 100 3
4, 9, 5
1.8865e-04 3.1003e-04 1.3642e-02 2.2184e-02
hiva 3075 1617 2
6, 5
1.2059e-03 3.2603e-03 7.5815e-03 1.5401e-02
imdb 3249 700 2
1, 1
1.7140e-05 1.7140e-05 1.0383e-02 1.0778e-02
secom 1566 590 2
2, 6
6.7170e-04 6.9461e-04 2.4978e-02 4.2688e-02
seismic 645 24 2
8, 1
9.2591e-03 1.1034e-03 1.8918e-02 2.2457e-02
sentiment 2747 200 2
13, 16
7.2360e-05 2.5743e-04 1.3689e-02 4.0064e-02
sms 834 500 2
2, 1
3.4828e-04 3.5066e-04 3.1853e-02 3.3163e-02
spambase 4600 57 2
15, 7
1.0266e-04 2.0145e-04 1.6376e-02 2.4708e-02
svhn 731 1024 10
6, 3, 3, 3, 3
3, 3, 2, 2, 2
1.1899e-02 5.7717e-03 1.2997e-02 1.4286e-02
@@ 5 Discussion and Conclusion
$=\exp\,\Biggl{[}-\frac{1}{2}\Biggl{(}\begin{bmatrix}y-P^{T}\mu\\ z-R^{T}\mu\end{bmatrix}^{T}U^{T}\Sigma^{-1}U\begin{bmatrix}y-P^{T}\mu\\ z-R^{T}\mu\end{bmatrix}\Biggr{)}\Biggr{]}\,$. (A.2)

### v2

@@ 17 Compute eigendecomposition of $\Sigma_{\mathrm{UA}}$
Dataset Size Dim. Labels GMM Components KL(PCA-KDE, wGMM-UAPCA) KL(PCA-KDE, UAPCA) SW2(PCA-KDE, wGMM-UAPCA) SW2(PCA-KDE, UAPCA)
bank 2058 63 2 16, 8 4.0804e-04 8.1839e-04 1.2428e-02 1.9343e-02
cifar10 3249 1024 10 1, 14, 1, 1, 1 1, 1, 13, 12 14 3.6982e-03 3.1093e-03 9.6107e-03 1.0517e-02
cnae9 1079 856 9 2, 1, 1, 1, 2 2, 1, 1, 1 4.3181e-04 4.4680e-04 8.9226e-03 8.8375e-03
coil20 1439 400 20 3, 3, 3, 3, 3 3, 3, 2, 3, 2 3, 2, 3, 3, 2 1, 2, 3, 3, 3 3.9615e-03 2.6747e-03 8.6938e-03 1.2066e-02
epileptic 5749 178 5 2, 3, 2, 1, 1 3.7940e-05 6.0040e-05 5.2632e-03 6.1744e-03
fashion_mnist 2999 784 10 1, 2, 12, 10, 11 2, 1, 2, 10, 2 8.4062e-03 9.9694e-03 8.8474e-03 1.2469e-02
fmd 996 1536 10 4, 4, 4, 4, 4 4, 5, 4, 3, 5 1.5755e-03 9.4136e-04 1.4576e-02 1.2633e-02
har 734 561 6 5, 5, 4, 4, 4, 5 3.0825e-03 2.2314e-03 8.5085e-03 1.1451e-02
hatespeech 3220 100 3 4, 9, 5 1.8865e-04 3.1003e-04 1.3642e-02 2.2184e-02
hiva 3075 1617 2 6, 5 1.2059e-03 3.2603e-03 7.5815e-03 1.5401e-02
imdb 3249 700 2 1, 1 1.7140e-05 1.7140e-05 1.0383e-02 1.0778e-02
secom 1566 590 2 2, 6 6.7170e-04 6.9461e-04 2.4978e-02 4.2688e-02
seismic 645 24 2 8, 1 9.2591e-03 1.1034e-03 1.8918e-02 2.2457e-02
sentiment 2747 200 2 13, 16 7.2360e-05 2.5743e-04 1.3689e-02 4.0064e-02
sms 834 500 2 2, 1 3.4828e-04 3.5066e-04 3.1853e-02 3.3163e-02
spambase 4600 57 2 15, 7 1.0266e-04 2.0145e-04 1.6376e-02 2.4708e-02
svhn 731 1024 10 6, 3, 3, 3, 3 3, 3, 2, 2, 2 1.1899e-02 5.7717e-03 1.2997e-02 1.4286e-02
@@ 5 Discussion and Conclusion
[…]